# Mehdi Zafari A. Lee Swindlehurst

### v1

@@ IV-B GNN Architecture
For prediction, dedicated MLP-based heads are used for each task: AP-user associations ($\mathbf{X}$) on communication edges, AP mode selection ($\boldsymbol{\tau}$) on AP nodes, target scheduling ($\mathbf{s}$) on target nodes, and AP-target links ($\mathbf{Y}^{\text{tx}}$, $\mathbf{Y}^{\text{rx}}$) on sensing edges. To jointly optimize all tasks, learnable log-variance parameters ($\log\sigma^{2}$) are introduced to adaptively weight each loss component, following the principle of task-uncertainty balancing. The overall loss combines binary cross-entropy terms for each output with additional regularization penalties enforcing AP RF chain limits and coupling consistency between variables (e.g., $\tau_{a}$ and $y_{a,t}^{\text{rx}}$). This formulation ensures stable multi-task learning and maintains consistency with the physical constraints of the original MILP problem. Once trained, the proposed ASSENT model efficiently infers near-optimal association and mode-selection decisions with sub-millisecond latency, offering a scalable and practical solution for real-time operation in distributed cell-free ISAC systems.
@@ V Evaluation
We first evaluate the MILP-based association optimization against four greedy baselines: MILP-aligned, channel-only, communication-only, and sensing-only. Here, “greedy” denotes myopic selection based on the maximum instantaneous link strength (product of priority weight and channel gain for each entity). For this evaluation, the priority weights $\lambda_{u}$ and $\lambda_{t}$ are set to one, and for each value of the trade-off parameter $\alpha\in[0,1]$, a total of $10,000$ independent channel realizations are generated. Fig. 1 (left) presents the empirical CDF of the achieved utilities for the MILP and greedy solutions, while Fig. 1 (right) illustrates how key network-level indicators, including fractions of UE coverage, scheduled-targets, transmit APs, and communication RF-chain share, evolve as $\alpha$ varies. These results confirm that the MILP formulation achieves the best utility and more balanced communication-sensing behavior than heuristic alternatives for various $\alpha$.
@@
Next we evaluate the ASSENT learning framework using the dataset generated from the MILP solutions. Unlike the fixed-parameter MILP evaluation, here $\alpha$, $\boldsymbol{\lambda}_{\text{cu}}$, and $\boldsymbol{\lambda}_{\text{tg}}$ are input variables and randomly sampled in the dataset generation process as discussed in Section IV-A. The three GNN variants NNConv, TransformerConv, and GATv2Con are implemented with hidden dimension $128$, three message-passing layers, batch size $16$, initial learning rate $10^{-3}$, cosine learning-rate scheduler, Adam optimizer, and early stopping. Fig. 2 (left) shows the training and validation loss trends with the learning-rate profile, and Fig. 2 (right) compares achieved objective values on the test set against the MILP benchmark. Table II summarizes detailed test-set performance metrics, including F1-score, precision, recall, and Brier score, along with true positive/negative (TP/TN) and false positive/negative (FP/FN) values for all five output variables. The high F1-scores (above 0.9 for all variables) demonstrate that ASSENT can accurately reproduce binary association and scheduling patterns learned from the MILP labels. Precision and recall remain well balanced, confirming that the model neither over-activates nor under-selects AP-entity links.
All codes, datasets, and trained checkpoints used to produce the above results are available in our open-source simulation repository [zafari2025assentrepo]. Overall, the experiments show that: (i) MILP yields consistent gains over greedy heuristics while revealing how network-level indicators trade off with $\alpha$, and (ii) ASSENT learns from MILP-optimal labels to deliver near-optimal utility and accurate decision variables on unseen test scenarios. Given that inference is a single forward pass, the learned models provide a practical surrogate for optimization, supporting real-time operation in distributed ISAC systems.
@@ VI Conclusion
In this paper we have presented ASSENT (learning-based ASSociation and ENTity selection), a GNN-based framework for efficient association optimization in distributed cell-free ISAC systems. We formulated the association problem as a MILP that captures interference, mode selection, and RF-chain constraints, and used its optimal solutions to supervise a heterogeneous GNN that infers association decisions from lightweight link statistics. Simulation results demonstrate that ASSENT achieves near-MILP utility while accurately reproducing the optimal association patterns. Once trained, a single forward pass efficiently replaces repeated MILP optimizations for real-time decision-making. All code and datasets have been released to support reproducibility and further research. Future work includes reinforcement learning for continual adaptation and experimental validation on distributed ISAC testbeds.

### v2

@@ IV-B GNN Architecture
For prediction, dedicated MLP-based heads are used for each task: AP-CU associations ($\mathbf{X}$) on communication edges, AP mode selection ($\boldsymbol{\tau}$) on AP nodes, target scheduling ($\mathbf{s}$) on target nodes, and AP-target links ($\mathbf{Y}^{\text{tx}}$, $\mathbf{Y}^{\text{rx}}$) on sensing edges. To jointly optimize all tasks, learnable log-variance parameters ($\log\sigma^{2}$) are introduced to adaptively weight each loss component, following the principle of task-uncertainty balancing. The overall loss combines binary cross-entropy terms for each output with additional regularization penalties enforcing AP RF chain limits and coupling consistency between variables (e.g., $\tau_{a}$ and $y_{a,t}^{\text{rx}}$). This formulation ensures stable multi-task learning and maintains consistency with the physical constraints of the original MILP problem. Once trained, the proposed ASSENT model efficiently infers near-optimal association and mode-selection decisions with sub-millisecond latency, offering a scalable and practical solution for real-time operation in distributed cell-free ISAC systems.
@@ V Evaluation
We first evaluate the MILP-based association optimization against four greedy baselines: MILP-aligned, channel-only, communication-only, and sensing-only. Here, “greedy” denotes myopic selection based on the maximum instantaneous link strength (product of priority weight and channel gain for each entity). For this evaluation, the priority weights $\lambda_{u}$ and $\lambda_{t}$ are set to one, and for each value of the trade-off parameter $\alpha\in[0,1]$, a total of $10,000$ independent channel realizations are generated. Fig. 1 (left) presents the empirical CDF of the achieved utilities for the MILP and greedy solutions, while Fig. 1 (right) illustrates how key network-level indicators, including fractions of UE coverage, scheduled-targets, transmit APs, and communication RF chain share, evolve as $\alpha$ varies. These results confirm that the MILP formulation achieves the best utility and more balanced communication-sensing behavior than heuristic alternatives for various $\alpha$.
@@
Next we evaluate the ASSENT learning framework using the dataset generated from the MILP solutions. Unlike the fixed-parameter MILP evaluation, here $\alpha$, $\boldsymbol{\lambda}_{\text{cu}}$, and $\boldsymbol{\lambda}_{\text{tg}}$ are input variables and randomly sampled in the dataset generation process as discussed in Section IV-A. The three GNN variants NNConv, TransformerConv, and GATv2Con are implemented with hidden dimension $H=128$, three message-passing layers, batch size $16$, initial learning rate $10^{-3}$, cosine learning-rate scheduler, Adam optimizer, and early stopping. Fig. 2 (left) shows the training and validation loss trends with the learning-rate profile, and Fig. 2 (right) compares achieved objective values on the test set against the MILP benchmark. Table II summarizes detailed test-set performance metrics, including F1-score, precision, recall, and Brier score, along with true positive/negative (TP/TN) and false positive/negative (FP/FN) values for all five output variables. The high F1-scores (above $0.84$ for all variables) demonstrate that ASSENT can accurately reproduce binary association and scheduling patterns learned from the MILP labels. Precision and recall remain well balanced, confirming that the model neither over-activates nor under-selects AP-entity links.
All codes, datasets, and trained checkpoints used to produce the above results are available in our open-source simulation repository [12]. Overall, the experiments show that: (i) MILP yields consistent gains over greedy heuristics while revealing how network-level indicators trade off with $\alpha$, and (ii) ASSENT learns from MILP-optimal labels to deliver near-optimal utility and accurate decision variables on unseen test scenarios. Given that inference is a single forward pass, the learned models provide a practical surrogate for optimization, supporting real-time operation in distributed ISAC systems.
@@ VI Conclusion
In this paper we have presented ASSENT (learning-based ASSociation and ENTity selection), a GNN-based framework for efficient association optimization in distributed cell-free ISAC systems. We formulated the association problem as a MILP that captures interference, mode selection, and RF chain constraints, and used its optimal solutions to supervise a heterogeneous GNN that infers association decisions from lightweight link statistics. Simulation results demonstrate that ASSENT achieves near-MILP utility while accurately reproducing the optimal association patterns. Once trained, a single forward pass efficiently replaces repeated MILP optimizations for real-time decision-making. All code and datasets have been released to support reproducibility and further research. Future work includes reinforcement learning for continual adaptation and experimental validation on distributed ISAC testbeds.